### v1

@@ Abstract.
We study the problem of broadcasting multiple messages in the CONGEST model. In this problem, a dedicated node $s$ possesses a set $M$ of messages with every message being of the size $O(\log n)$ where $n$ is the total number of nodes. The objective is to ensure that every node in the network learns all messages in $M$. The execution of an algorithm progresses in rounds and we focus on optimizing the round complexity of broadcasting multiple messages.
@@
Our primary contribution is a randomized algorithm designed for networks modeled as random graphs. The algorithm succeeds with high probability and achieves round complexity that is optimal up to a polylogarithmic factor. It leverages a multi-COBRA primitive, which uses multiple branching random walks running in parallel. To the best of our knowledge, this approach has not been applied in distributed algorithms before. A crucial aspect of our method is the use of these branching random walks to construct an optimal (up to a polylogarithmic factor) tree packing of a random graph, which is then used for efficient broadcasting. This result is of independent interest.
@@
We also prove the problem to be NP-hard in a centralized setting and provide insights into why straightforward lower bounds, namely graph diameter and $\frac{|M|}{minCut}$, can not be tight.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Network topologies of major distributed systems, such as Bitcoin (nakamoto2008bitcoin, ), Ethereum (buterin2014next, ), and Torrent (cohen2003incentives, ), are often designed to emulate the properties of random graphs. Empirical studies further confirm that real-world networks frequently exhibit characteristics akin to random graphs (gao2019topologyEth, ; delgado2019txprobeBTC, ; dale2008evolutionTorrent, ). This observation has naturally motivated a significant body of research to focus on algorithms and protocols tailored for random graph structures.
@@ 1. Introduction
One of the fundamental communication tasks in such distributed systems is the efficient broadcast of large messages. In blockchains, this could be the dissemination of a block of transactions, while in peer-to-peer networks like torrents, it may involve distributing large files. The design of robust and efficient broadcast algorithms is critical for maintaining the performance and scalability of these systems.
@@
In this work, we present an algorithm that addresses the problem of broadcasting large messages in distributed systems with a random graph topology. Our solution achieves near-optimal broadcast time while imposing no requirements on the underlying network structure beyond its random properties. This generality ensures that our approach is applicable to a wide range of real-world systems, including various blockchain architectures, and highlights its potential as a versatile tool for scalable and efficient information dissemination.
1.1. Model and Problem
@@
The CONGEST model (peleg2000distributed, ) is defined as follows. The network is modeled as a graph with $n$ nodes, where execution progresses in synchronous rounds. In each round, a node can send a message of size $O(\log n)$ bits to each of its neighbors. Nodes do not have prior knowledge of the network topology but are assumed to have unique identifiers that fit within $O(\log n)$ bits.
@@
Although the CONGEST model has been extensively studied over the past two decades, the fundamental problem of broadcasting multiple messages remains incompletely understood.
Definition 0 (Multi-message broadcast).
@@
A dedicated node $s$ possesses a set $M$ of messages, where each message $m\in M$ has a size of $O(\log n)$ bits. The objective is to ensure that every node in the network learns all messages in $M$.
1.2. On the Universal Optimality
@@
As pointed out by Ghaffari (ghaffari2015distributed, ), the problem suggests an $\Omega(D+k)$ round complexity lower bound, where $D$ is the diameter of the graph and $k$ is the number of messages $|M|$. For example, consider a path graph with $s$ as its first node. Any algorithm would require at least $D+k-1$ rounds to transmit all messages to the last node. However, this bound is existential, meaning there exists a graph for which $\Omega(D+k)$ rounds are needed. In contrast, consider a complete graph with $k=n$. Here, broadcasting can be completed in 2 rounds, which is significantly better than the $\Omega(k)=\Omega(n)$ bound suggested by the path graph example. This paper presents an algorithm that achieves universal optimality (garay1998sublinear, ) on random graphs. Specifically, for a random graph $G$, the algorithm completes the multi-message broadcast in $\tilde{O}(OPT(G))$ rounds with high probability, where $OPT(G)$ denotes the best possible round complexity for $G$ and $\tilde{O}$ hides $polylog(k,n)$ factors.
1.3. Previous Work
The first work to address universal optimality for the multi-message broadcast problem in the CONGEST model was “Distributed Broadcast Revisited: Towards Universal Optimality“ by Ghaffari (ghaffari2015distributed, ). In that paper, the algorithm consists of two phases: (1) constructing a tree packing, and (2) performing the broadcast using the constructed tree packing. A tree packing of a graph $G$ is a collection of spanning subtrees of $G$. The tree packing is characterized by three parameters: (1) its size $S$, i.e., the number of trees, (2) its height $H$, i.e., the maximal height of a tree, and (3) its weight $W$, i.e., the maximal number of trees sharing a single edge. With a tree packing, one can complete a multi-message broadcast in $O((H+\frac{k}{S})\cdot W)$ rounds by splitting messages uniformly across the trees and propagating them sequentially within each tree. However, the limitation of (ghaffari2015distributed, ) is that constructing the tree packing requires $\tilde{\Omega}(D+k)$ rounds, preventing the approach from achieving universal optimality.
A subsequent work, “Fast Broadcast in Highly Connected Networks“ [preprint 2024] by Ghaffari et al. (chandra2024fast, ), considered the tree packing approach on highly connected graphs, i.e., graphs with high edge connectivity $\lambda$. The primary result of this work is an algorithm that runs in $\tilde{O}(\frac{n+k}{\lambda})$ rounds. This complexity is optimal when $k=\Omega(n)$, as $\frac{k}{\lambda}$ represents an information-theoretic lower bound. However, the algorithm may incur a $\tilde{\Omega}(n)$ factor overhead in cases where $\lambda$ and $k$ are small compared to $n$.
Notably, both (ghaffari2015distributed, ) and (chandra2024fast, ) consider a slightly more general problem where initially $M$ is not necessarily known to a single node but different nodes can possess different parts of it. We adhere to our version, where $M$ is initially held by a single node, as it simplifies the presentation. Importantly, when a tree packing is available, the multiple-source version can be reduced to the single-source version without increasing the round complexity (see Remark 14).
1.4. Preliminaries and Notation
@@
An Erdős–Rényi graph $G(n,p)$is a graph on $n$ vertices where each edge exists independently from others with probability $p$ (erd6s1960evolution, ). Throughout the paper, for a graph $G$, $E(G)$ is the edge set, $V(G)$ is the vertex set, $D(G)$ is the diameter, $\delta(G)$ is the smallest vertex degree, and $\Delta(G)$ is the largest vertex degree. We do not explicitly specify $G$ if it is obvious from the context, e.g., we can write $\delta$ instead of $\delta(G)$. With high probability (w.h.p.) means with a probability of at least $1-O(\frac{1}{n^{C}})$ for some constant $C>0$, with the probability being taken over both the randomness of the graph and the random bits of the algorithm. We assume that $\tilde{O}$ and $\tilde{\Omega}$ hide $polylog(k,n)$ factors.
1.5. Our Contribution
In the present paper, we provide an algorithm that is universally optimal w.h.p. when the network is modeled as an Erdős–Rényi graph $G(n,p)$.
Theorem 2.
@@ 1.5. Our Contribution
Throughout the paper, let $C_{p}$ denote a sufficiently large constant. [^1] It suffices to take $C_{p}=2700$. For an Erdős–Rényi graph $G(n,p)$ with $p\geq\frac{C_{p}\log n}{n}$, there exists a distributed randomized algorithm that completes the broadcast in $\tilde{O}(D(G)+\frac{k}{\delta(G)})$ rounds w.h.p.
Remark 3.
The round complexity of $\tilde{O}(D(G)+\frac{k}{\delta(G)})$ is optimal up to a polylogarithmic factor since $D(G)$ and $\frac{k}{\delta(G)}$ are both straightforward lower bounds.
Remark 4.
@@
The condition $p=\Omega(\frac{\log n}{n})$ is necessary, since for $p\leq\frac{\log n}{n}$, there is a constant probability that the graph is disconnected (erd6s1960evolution, ).
@@
To obtain Theorem 2, we use the following result of independent interest:
Theorem 5.
@@
For an Erdős–Rényi graph $G(n,p)$ with $p\geq\frac{C_{p}\log n}{n}$, there exists a distributed randomized algorithm that produces a tree packing of size $\delta(G)$, height $\tilde{O}(D(G))$ and weight $O(\sqrt{\log n})$ w.h.p.
@@
We construct the latter algorithm by utilizing multiple Coalescing Branching Random Walks (cooper2016coalescing, ) that run in parallel. To the best of our knowledge, this technique has never been used before in the context of distributed algorithms.
@@
Finally, to map the terrain of the problem, we prove its NP-hardness in the centralized case. Also, we provide an instance of a problem where $D(G)$ and $\frac{k}{minCut(G)}$ two straightforward lower bounds for round complexity - are both $O(1)$, while the optimal round complexity is $\Omega(\sqrt{k})$.
2. Related Work
Tree Packing. The problem of tree packing has been extensively studied, as summarized in the survey by Palmer (palmer2001spanning, ). Foundational results in this area include those by Tutte (tutte1961problem, ) and Nash-Williams (nash1961edge, ), who demonstrated that an undirected graph with edge connectivity $\lambda$ contains a tree packing of size $\lfloor\frac{\lambda}{2}\rfloor$. Edmond (edmonds1972, ) extended this result to directed graphs, showing that such graphs always contain $\lambda$ pairwise edge-disjoint spanning trees rooted at a sender $s\in V$, where $\lambda$ is the minimum number of edges that must be removed to make some node unreachable from $s$. However, these results do not address the height of the tree packing.
Chuzhoy et al. (chuzhoy2020packing, ) tackled the challenge of finding tree packings with small height. They presented a randomized algorithm that, given an undirected graph with edge connectivity $\lambda$ and diameter $D$, outputs with high probability a tree packing of size $\lfloor\frac{\lambda}{2}\rfloor$, weight $2$, and height $O((101k\log n)^{D})$.
Tree packing on random graphs was studied by Gao et al. (gao2014arboricity, ), who showed that asymptotically almost surely, the size of a spanning tree packing of weight one for a Erdős–Rényi graph $G(n,p)$ is $\min\left\{\delta(G),\frac{|E(G)|}{n-1}\right\}$, which corresponds to two straightforward upper bounds.
In the CONGEST model, tree packing was investigated by Censor-Hillel et al. (censor2014distributed, ). They proposed an algorithm to decompose an undirected graph with edge connectivity $\lambda$ into fractionally edge-disjoint weighted spanning trees with total weight $\lceil\frac{\lambda-1}{2}\rceil$ in $\tilde{O}(D+\sqrt{n\lambda})$ rounds. Furthermore, they proved a lower bound of $\tilde{\Omega}(D+\sqrt{\frac{n}{\lambda}})$ on the number of rounds required for such a decomposition.
@@ 2. Related Work
The key difference between the network information flow problem and the multi-message broadcast in CONGEST is that in our problem, the focus is on round complexity, whereas in the information flow problem, the solution is a ”static” assignment of messages to edges.
@@
Branching Random Walks in Networks. The cover time of a random walk (lawler2010random, ) on a graph is the time needed for a walk to visit each node at least once. Unfortunately, the expected value of this quantity is $\Omega(n\log n)$ even for a clique, making this primitive less useful in practical applications. Consequently, several attempts have been made to accelerate the cover time. Alon et al. (alon2008many, ) proposed initiating multiple random walks from a single source. Subsequent work by Elsässer and Sauerwald refined their bounds, demonstrating that $r$ random walks can yield a speed-up of $r$ times for many graph classes. Variations of multiple random walks have been applied in the CONGEST model to approximate the mixing time (molla2017distributed, ), perform leader election (kutten2015sublinear, ; gilbert2018leader, ), and evaluate network conductance (fichtenberger2018two, ; batu2024all, ).
@@
A branching random walk (shi2015branching, ) (BRW) modifies the classical random walk by allowing nodes to emit multiple copies of a walk upon receipt, rather than simply relaying it. This branching behavior potentially leads to exponential growth in the number of walks traversing the graph, significantly reducing the cover time. Roche (roche2017robust, ), in his Ph.D. thesis “Robust Local Algorithms for Communication and Stability in Distributed Networks“ [2017], utilized BRWs to maintain the expander topology of a network despite adversarial node deletions and insertions. Gerraoui et al. (guerraoui2023inherent, ), in “On the Inherent Anonymity of Gossiping“, demonstrated that BRWs can enhance privacy by obscuring the source of gossip within a network. Recently, Aradhya et al. (aradhya_et_al:LIPIcs.OPODIS.2024.36, ) in “Distributed Branching Random Walks and Their Applications“ employed BRWs to address permutation routing problems on subnetworks in the CONGEST model.

### v2

@@ Abstract
We study the problem of broadcasting multiple messages in the CONGEST model. In this problem, a dedicated source node $s$ possesses a set $M$ of messages with every message of size $O(\log n)$ where $n$ is the total number of nodes. The objective is to ensure that every node in the network learns all messages in $M$. The execution of an algorithm progresses in rounds, and we focus on optimizing the round complexity of broadcasting multiple messages.
@@
Our primary contribution is a randomized algorithm for networks with expander topology, which are widely used in practice for building scalable and robust distributed systems. The algorithm succeeds with high probability and achieves a round complexity that is optimal up to a factor of the network’s mixing time and polylogarithmic terms. It leverages a multi-COBRA primitive, which uses multiple branching random walks running in parallel. To the best of our knowledge, this approach has not been applied in distributed algorithms before. A crucial aspect of our method is the use of these branching random walks to construct an optimal (up to a polylogarithmic factor) tree packing of a random graph, which is then used for efficient broadcasting. This result is of independent interest.
@@
We also prove the problem to be NP-hard in a centralized setting and provide insights into why straightforward lower bounds for general graphs, namely graph diameter and $\frac{|M|}{\textit{minCut}}$, cannot be tight.
keywords:
Distributed algorithms, Expander Graphs, Random graphs, Broadcasting, Branching random walks, Tree packing, CONGEST model.
1 Introduction
Expander graphs possess a unique combination of highly desirable properties for networked systems, including low diameter, high connectivity, rapid mixing, and strong resilience to node failures. These characteristics make them ideal topologies for ensuring fast information dissemination, robust communication, and efficient resource discovery in distributed environments. As a result, overlay network topologies of major distributed systems are often designed to emulate the properties of expander graphs [28, 43, 34, 50, 14, 4, 51, 46]. This practical adoption of high-expansion structures further reinforces the motivation to study and design algorithms specifically optimized for expander graphs, aiming to leverage their theoretical advantages in real-world distributed systems.
@@ 1 Introduction
One of the fundamental communication tasks in such distributed systems is to efficiently broadcast multiple messages. In blockchains, this could be the dissemination of transactions, while in peer-to-peer networks such as Torrent, it may involve distributing chunks of a large file. The design of robust and efficient broadcast algorithms is critical for maintaining the performance and scalability of these systems.
@@
In this work, we present an algorithm that addresses the problem of broadcasting multiple messages in distributed systems with an expander graph topology. Our approach broadcasts information nearly optimally, with an overhead proportional to the network’s mixing time and a small polylogarithmic factor while imposing no requirements on the underlying network topology beyond its expansion properties. This generality ensures that our approach is applicable to a wide range of real-world systems and highlights its potential as a versatile tool for scalable and efficient information dissemination. Let us now formally define the setting.
1.1 Model and Problem
@@
The CONGEST model [47] is defined as follows. The network is modeled as a graph with $n$ nodes, where execution progresses in synchronous rounds. In each round, a node can send a message of size $O(\log n)$ bits to each of its neighbors. Nodes do not have prior knowledge of the network topology but are assumed to have unique identifiers that fit within $O(\log n)$ bits.
@@
Although the CONGEST model has been extensively studied over the past two decades, the fundamental problem of broadcasting multiple messages remains unsolved for general topologies.
Definition 1.1 (Multi-message broadcast).
@@
A dedicated source node $s$ possesses a set $M$ of messages, where each message $m\in M$ has a size of $O(\log n)$ bits. The objective is to ensure that every node in the network learns all messages in $M$.
@@
As pointed out by Ghaffari [23], the problem suggests an $\Omega(D+k)$ round complexity lower bound, where $D$ is the diameter of the graph and $k$ is the number of messages $|M|$. For example, consider a path graph with $s$ as its first node. Any algorithm would require at least $D+k-1$ rounds to transmit all messages to the last node. However, this bound is only existential, meaning there exists a graph for which $\Omega(D+k)$ rounds are needed. In contrast, consider a complete graph with $k=n$. Here, broadcasting can be completed in $2$ rounds, which is significantly better than the $\Omega(k)=\Omega(n)$ bound suggested by the path graph example. Our paper presents an algorithm that achieves universal optimality [22] on expander graphs. Specifically, for any expander graph, our algorithm completes the multi-message broadcast in a number of rounds that is within a small overhead of the best possible for that specific graph. Before we state our results formally, we need to introduce helpful terminology.
1.2 Preliminaries and Notation
@@
Throughout the paper, for a graph $G$, $E(G)$ is the edge set, $V(G)$ is the vertex set, $D(G)$ is the diameter, $d_{G}(v)$ is the degree of a node $v$ in $G$, $\delta(G)$ is the smallest vertex degree, and $\Delta(G)$ is the largest vertex degree. We do not explicitly specify $G$ if it is obvious from the context, e.g., we can write $\delta$ instead of $\delta(G)$. With high probability (w.h.p.) means with a probability of at least $1-O(\frac{1}{n^{C}})$ for some constant $C>0$, with the probability being taken over both the randomness of the graph (when we assume random graphs) and the random bits of the algorithm. We assume that $\tilde{O}$ and $\tilde{\Omega}$ hide $polylog(k,n)$ factors.
In multiple places in this paper, we are using classical Chernoff bounds. A reader can find formal statements in Appendix A.
One of the key graph-theoretical components in our approach is tree packing. A tree packing of a graph $G$ is a collection of spanning subtrees of $G$. The tree packing is characterized by three parameters: (1) its size $S$, i.e., the number of trees, (2) its height $H$, i.e., the maximal height of a tree, and (3) its weight $W$, i.e., the maximal number of trees sharing a single edge.
The present work focuses specifically on two graph families, namely Erdős–Rényi graphs and expanders. An Erdős–Rényi graph $G(n,p)$ is a graph on $n$ vertices where each edge exists independently from others with probability $p$ [18]. We refer to a graph as an expander if it has small (polylogarithmic) mixing time [^1] An alternative way is to say that expander graphs are those that feature an inverse polylogarithmic conductance; these two notions are equivalent. See Appendix A for a definition of conductance..
For our purposes, it will be convenient to define the mixing time of an undirected graph by reconsidering it as being bidirected, that is, with each undirected edge $(u,v)$ replaced with $(u,v)$ and $(v,u)$. The mixing time $\tau_{mix}$ of a bidirected graph is defined as follows. Consider a lazy random process that starts at an arbitrary edge of the graph. At each step, with probability $\frac{1}{2}$, the process remains at the current edge, and with probability $\frac{1}{2}$, it transitions to a uniformly random adjacent edge (a directed edge $e_{2}$ is adjacent to a directed edge $e_{1}$ if they are of the form $e_{1}=(u,v)$ and $e_{2}=(v,w)$). It is known that this process admits a stationary distribution $\pi$, which is uniform over all edges. Moreover, regardless of the starting edge, the distribution $D_{t}$ of the walk after $t$ steps converges to $\pi$. We define $\tau_{mix}$ as the smallest $t$ for which $D_{t}$ is inverse-polynomially close to $\pi$. For a formal definition, please refer to Appendix A. We are now ready to formally state our results.
1.3 Our Contribution
Our main result is an algorithm to solve the multi-message broadcast problem with only an overhead of $\tilde{O}(\tau_{mix})$.
Theorem 1.2.
There exists a randomized distributed algorithm that for any graph $G$ solves the multi-message broadcast problem in $O(\log^{3}n\cdot\tau_{mix}\cdot\mathrm{OPT})$ rounds with high probability, where $\tau_{mix}$ is the mixing time of the network graph and $\mathrm{OPT}$ is the optimal round complexity for the given problem instance.
We design the algorithm for a general expander by building on the algorithm for random graphs, which achieves near-optimal performance in an important special case when the network is modeled as an Erdős–Rényi graph $G(n,p)$.
Throughout the paper, let $C_{p}$ denote a sufficiently large constant. [^2] It suffices to take $C_{p}=2700$. It was not a concern for the present work to optimize this constant.
Theorem 1.3.
@@ 1.3 Our Contribution
For an Erdős–Rényi graph $G(n,p)$ with $p\geq\frac{C_{p}\log n}{n}$, there exists a distributed randomized algorithm that completes the broadcast in $O(\log^{2}n+\log n\cdot\frac{k}{\delta(G)})$ rounds w.h.p.
@@
Note that the round complexity of $O(\log^{2}n+\log n\cdot\frac{k}{\delta(G)})$ is optimal up to an additive $\log^{2}n$ term and a multiplicative $\log n$ factor since $\frac{k}{\delta(G)}$ is a straightforward lower bound. Also, the condition $p=\Omega(\frac{\log n}{n})$ is necessary, since for $p\leq\frac{\log n}{n}$, there is a constant probability that the graph is disconnected [18].
@@
To obtain Theorem 1.3, we use the following result of independent interest, which contributes to a line of work [24, 11, 23, 7, 21] on low-diameter tree packing:
Theorem 1.4.
@@
For an Erdős–Rényi graph $G(n,p)$ with $p\geq\frac{C_{p}\log n}{n}$, there exists a distributed randomized algorithm that produces a tree packing of size $\delta(G)$, height $O(\log n)$ and weight $O(\log n)$ w.h.p.
@@
We construct the latter algorithm by utilizing multiple Coalescing Branching Random Walks [12] that run in parallel. To the best of our knowledge, this technique has never been used before in the context of distributed algorithms.
@@
Finally, to map the terrain of the problem, we prove that it is NP-hard in the centralized case to achieve the optimal number of rounds for general graphs. Also, we provide an instance of a problem where $D(G)$ and $\frac{k}{\textit{minCut}(G)}$ – two straightforward lower bounds for round complexity – are both $O(1)$, while the optimal round complexity is $\Omega(\sqrt{k})$.
2 Related Work
Previous Work. The first work to address universal optimality for the multi-message broadcast problem in the CONGEST model was by Ghaffari [23]. In that paper, the algorithm consists of two phases: (1) constructing a tree packing, and (2) performing the broadcast using the constructed tree packing. With a tree packing of height $H$, size $S$ and weight $W$, one can complete a multi-message broadcast in $O((H+\frac{k}{S})\cdot W)$ rounds by splitting messages uniformly across the trees and propagating them sequentially within each tree. However, the limitation of [23] is that constructing the tree packing requires $\tilde{\Omega}(D+k)$ rounds, preventing the approach from achieving universal optimality.
A subsequent work by Ghaffari et al. [8] considered the tree packing approach on highly connected graphs, i.e., graphs with high edge connectivity $\lambda$. The primary result of this work is an algorithm that runs in $\tilde{O}(\frac{n+k}{\lambda})$ rounds. This complexity is optimal when $k=\Omega(n)$, as $\frac{k}{\lambda}$ represents an information-theoretic lower bound. However, the algorithm may incur a $\tilde{\Omega}(n)$ factor overhead in cases where $\lambda$ and $k$ are small compared to $n$.
Notably, both [23] and [8] consider a slightly more general problem where initially $M$ is not necessarily known to a single node, but different nodes can possess disjoint subsets of $M$. We adhere to our version, where $M$ is initially held by a single node, as it simplifies the presentation. Importantly, when a tree packing is available, the multiple-source version can be reduced to the single-source version without increasing the round complexity (see Remark 4.18).
Routing. One fundamental information-dissemination problem in distributed computing is routing, where the goal is to deliver a set of messages from source nodes to their respective destination nodes. Unlike broadcast, which involves sending the same message(s) to all nodes, routing requires sending individual message(s) for each source-destination pair.
Ghaffari et al. [25] approach this problem for expanders by constructing a hierarchy of recursively embedded Erdős–Rényi graphs, and achieve routing in $\tau_{\text{mix}}\cdot 2^{O(\sqrt{\log n\log\log n})}$ rounds. Note that $2^{O(\sqrt{\log n\log\log n})}$ dominates $\log^{c}n$ for any constant $c$. This result was further improved in [26], which reduces the round complexity to $O(2^{\sqrt{\log n}})$. Subsequently, Chang and Thatchaphol [9] presented a deterministic version of expander routing, matching (up to polylogarithmic factors) the round complexity of the randomized algorithm by Ghaffari et al.
For general graphs, Haeupler et al. [30] provide a routing algorithm that runs in $\mathrm{poly}(D)\cdot n^{o(1)}$ rounds. Their approach leverages expander decomposition and hop-constrained expanders—subgraphs with small diameter and strong expansion properties. In fact, [30] obtain a stronger result: given that for every source-sink pair $(s_{i},t_{i})$, the source $s_{i}$ is at most $h$ hops from its destination $t_{i}$, routing can be completed in $O(D+\mathrm{poly}(h))\cdot n^{o(1)}$ rounds.
@@ 2 Related Work
The key difference between the network information flow problem and the multi-message broadcast in CONGEST is that in our problem, the focus is on round complexity, whereas in the information flow problem, the solution is a "static" assignment of messages to edges, and the focus is on throughput.
Broadcast in the Presence of Churn. To account for the instability of connections in distributed systems, researchers often consider models where the graph topology undergoes modifications while the algorithm is running. Kuhn et al. [37] introduced a model where, in each round, the communication graph is chosen adversarially, and achieved a round complexity of $O(nk)$ for broadcasting $k$ messages in this setting. As a follow-up, Dutta et al. [16] established a lower bound of $\tilde{\Omega}(nk)$ for this problem, providing a precise sequence of adversarially constructed graphs that makes the problem difficult. However, Dinitz et al. [13] showed that this lower bound is “fragile.” Specifically, they demonstrated that the hard case is carefully constructed to prevent algorithms from succeeding. They did so by introducing the concept of $l$-smoothing to the adversarial sequence of graphs, where each graph in the sequence additionally gets $l$ edges added at random. Surprisingly, even with $l=1$, an algorithm that broadcasts a random message at possession to its neighbors in each round achieves a round complexity of $\tilde{O}(n+k^{3})$ w.h.p., effectively breaking the lower bound of [16].
Tree Packing. The problem of tree packing has been extensively studied, as summarized in the survey by Palmer [45]. Foundational results in this area include those by Tutte [53] and Nash-Williams [44], who demonstrated that an undirected graph with edge connectivity $\lambda$ contains a tree packing of size $\lfloor\frac{\lambda}{2}\rfloor$. Edmond [17] extended this result to directed graphs, showing that such graphs always contain $\lambda$ pairwise edge-disjoint spanning trees rooted at a sender $s\in V$, where $\lambda$ is the minimum number of edges that must be removed to make some node unreachable from $s$. However, these results do not address the height of the tree packing.
Chuzhoy et al. [11] tackled the challenge of finding tree packings with small height. They presented a randomized algorithm that, given an undirected graph with edge connectivity $\lambda$ and diameter $D$, outputs with high probability a tree packing of size $\lfloor\frac{\lambda}{2}\rfloor$, weight $2$, and height $O((101k\log n)^{D})$.
Tree packing on random graphs was studied by Gao et al. [21], who showed that asymptotically almost surely, the size of a spanning tree packing of weight one for a Erdős–Rényi graph $G(n,p)$ is $\min\left\{\delta(G),\frac{|E(G)|}{n-1}\right\}$, which corresponds to two straightforward upper bounds.
In the CONGEST model, tree packing was investigated by Censor-Hillel et al. [7]. They proposed an algorithm to decompose an undirected graph with edge connectivity $\lambda$ into fractionally edge-disjoint weighted spanning trees with total weight $\lceil\frac{\lambda-1}{2}\rceil$ in $\tilde{O}(D+\sqrt{n\lambda})$ rounds. Furthermore, they proved a lower bound of $\tilde{\Omega}(D+\sqrt{\frac{n}{\lambda}})$ on the number of rounds required for such a decomposition.
@@
Branching Random Walks in Networks. The cover time of a random walk [39] on a graph is the time needed for a walk to visit each node at least once. Unfortunately, the expected value of this quantity is $\Omega(n\log n)$ even for a clique, making this primitive less useful in practical applications. Consequently, several attempts have been made to accelerate the cover time. Alon et al. [2] proposed initiating multiple random walks from a single source. Subsequent work by Elsässer and Sauerwald refined their bounds, demonstrating that $r$ random walks can yield a speed-up of $r$ times for many graph classes. Variations of multiple random walks have been applied in the CONGEST model to approximate the mixing time [42], perform leader election [38, 27], and evaluate network conductance [19, 5].
@@
A branching random walk [49] (BRW) modifies the classical random walk by allowing nodes to emit multiple copies of a walk upon receipt, rather than simply relaying it. This branching behavior potentially leads to exponential growth in the number of walks traversing the graph, significantly reducing the cover time. Gerraoui et al. [29] demonstrated that BRWs can enhance privacy by obscuring the source of gossip within a network. Recently, Aradhya et al. [3] employed BRWs to address permutation routing problems on subnetworks in the CONGEST model.